OpenStudy (hellomynameiskam):

Simplify. 16 – (–12) ÷ (–2) A. –2 B. 22 C. 10 D. –14

OpenStudy (gabebae):

Use PEMDAS

OpenStudy (alexandervonhumboldt2):

-/-=+. 16-(-12)/(-2)=16-12/2=16-6=???

OpenStudy (anonymous):

Remember PEMDAS, parentheses, exponents, division and multiplication, addition subtraction

OpenStudy (alexandervonhumboldt2):

first divide, then subtarct

OpenStudy (gabebae):

Yup

OpenStudy (alexandervonhumboldt2):

use PEDMAS

OpenStudy (hellomynameiskam):

so it's C

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (hellomynameiskam):

thanks

OpenStudy (alexandervonhumboldt2):

np